Find the equation of the curve whose slope at any point is 𝑑𝑦/𝑑π‘₯ = π‘₯^2 √π‘₯ and which passes through the point (1,0).


1
Expert's answer
2021-02-28T16:36:17-0500

The slope of the curve is given by "\frac{dy}{dx}=x^2.\sqrt{x}" .....(1)

Integrating both side of (1) with respect to "x" ,we get

"\intop dy=\intop x^2.\sqrt{x} dx"

"\implies \intop dy=\intop x^{\frac{5}{2}} dx"

"\implies y=\frac{x^{(\frac{5}{2}+1)}}{\frac{5}{2}+1}+C"

"\implies y=\frac{2}{7}.x^{\frac{7}{2}}+C" [where "C" is an integrating constant]

As the curve passes through the point "(1,0)" we have,

"0=\frac{2}{7}.1^{\frac{7}{2}}+C"

"\implies C=-\frac{2}{7}"

Therefore the required equation of the curve is "y=\frac{2}{7}.x^{\frac{7}{2}}-\frac{2}{7}"